# FlashSinkhorn: IO-Aware Entropic Optimal Transport

### v1

@@ Abstract
Entropic optimal transport (EOT) via Sinkhorn iterations is widely used in modern machine learning, yet GPU solvers remain inefficient at scale. Tensorized implementations suffer quadratic HBM traffic from dense $n\times m$ interactions, while existing online backends avoid storing dense matrices but still rely on generic tiled map-reduce reduction kernels with limited fusion. We present FlashSinkhorn, an IO-aware EOT solver for squared Euclidean cost that rewrites stabilized log-domain Sinkhorn updates as row-wise LogSumExp reductions of biased dot-product scores, the same normalization as transformer attention. This enables FlashAttention-style fusion and tiling: fused Triton kernels stream tiles through on-chip SRAM and update dual potentials in a single pass, substantially reducing HBM IO per iteration while retaining linear-memory operations. We further provide streaming kernels for transport application, enabling scalable first- and second-order optimization. On A100 GPUs, FlashSinkhorn achieves up to $32\times$ forward-pass and $161\times$ end-to-end speedups over state-of-the-art online baselines on point-cloud OT, improves scalability on OT-based downstream tasks. For reproducibility, we release an open-source implementation at https://github.com/ot-triton-lab/ot_triton.
@@ 2.1 EOT and Sinkhorn Algorithm
where $\Pi(\mathbf{a},\mathbf{b})=\{P\in\mathbb{R}_{\geq 0}^{n\times m}:P\mathbbm{1}_{m}=\mathbf{a},P^{\top}\mathbbm{1}_{n}=\mathbf{b}\}$ is the transport polytope, $C\in\mathbb{R}^{n\times m}$ is the cost matrix with entries $C_{ij}=\|\mathbf{x}_{i}-\mathbf{y}_{j}\|_{2}^{2}$, i.e., the squared Euclidean distance, and $\textrm{KL}\left(P\middle\|\mathbf{a}\otimes\mathbf{b}\right):=\sum_{i=1}^{n}\sum_{j=1}^{m}P_{ij}\Big(\log\frac{P_{ij}}{a_{i}b_{j}}\Big)$, denotes the relative entropy with $\varepsilon>0$ the regularization strength.
@@ 3.1 Streaming Sinkhorn Iterations
5:  Load $Q_{I}$ to on-chip SRAM
@@
8:   Load $K_{J}$, $\hat{\mathbf{g}}_{J}$ and $\bm{\delta}_{J}$ to on-chip SRAM
9:   $S\leftarrow(Q_{I}K_{J}^{\top}+\hat{\mathbf{g}}_{J}+\bm{\delta}_{J})/\varepsilon$ // score tile + bias
@@ 3.2 Streaming Transport Matrix Application
5:  Load $Q_{I}$, $\hat{\mathbf{f}}_{I}$, $\mathbf{a}_{I}$ to on-chip SRAM
6:  $\mathbf{m}_{I}\leftarrow$-$\infty$, $\mathbf{s}_{I}\leftarrow 0$, $O_{I}\leftarrow 0$
@@
8:   Load $K_{J}$, $\hat{\mathbf{g}}_{J}$, $\bm{\delta}_{J}$ and $V_{J}$ to on-chip SRAM
9:   $S\leftarrow(Q_{I}K_{J}^{\top}+\hat{\mathbf{g}}_{J}+\bm{\delta}_{J})/\varepsilon$ // score tile + bias
@@
13:   $O_{I}\leftarrow e^{\mathbf{m}_{I}-\mathbf{m}_{\mathrm{new}}}\odot O_{I}+e^{S-\mathbf{m}_{\mathrm{new}}}\,V_{J}$
14:   $\mathbf{m}_{I}\leftarrow\mathbf{m}_{\mathrm{new}}$
15:  end for
16:  $\hat{\mathbf{f}}^{+}_{I}\leftarrow-\varepsilon\cdot\big(\mathbf{m}_{I}+\log\mathbf{s}_{I}\big)$ // one-step (10)
17:  $\mathbf{r}_{I}\leftarrow\mathbf{a}_{I}\odot\exp\!\left((\hat{\mathbf{f}}_{I}-\hat{\mathbf{f}}^{+}_{I})/\varepsilon\right)$ // $\mathbf{r}_{I}=(P\mathbbm{1})_{I}$
18:  $\mathrm{out}_{I}\leftarrow\mathrm{diag}(\mathbf{r}_{I})\,\mathrm{diag}(\mathbf{s}_{I})^{-1}\,O_{I}$
19:  Write $\mathrm{out}_{I}$ to HBM
20:end for
@@ Appendix B Derivation of Sinkhorn algorithm
where $\Pi(\mathbf{a},\mathbf{b})=\{P\in\mathbb{R}_{\geq 0}^{n\times m}:P\mathbbm{1}_{m}=\mathbf{a},P^{\top}\mathbbm{1}_{n}=\mathbf{b}\}$ , $C\in\mathbb{R}^{n\times m}$ is the cost matrix with entries $c_{ij}=\|\mathbf{x}_{i}-\mathbf{y}_{j}\|^{2}_{2}$, and $\mathrm{KL}(P\|\mathbf{a}\otimes\mathbf{b})=\sum_{ij}P_{ij}\log\left(\frac{P_{ij}}{a_{i}b_{j}}\right)$ is the Kullback–Leibler (KL) divergence (also called relative entropy) and $\varepsilon$ is the entropy regularization strength.
@@
$\mathcal{L}$ $({\mathbf{f}},{\mathbf{g}},P)=\sum_{ij}\Big[c_{ij}P_{ij}+\varepsilon P_{ij}\log\left(\frac{P_{ij}}{a_{i}b_{j}}\right)-f_{i}(P_{ij}-a_{i})-g_{j}(P_{ij}-b_{j})\Big].$
where $\mathbf{f}$ and $\mathbf{g}$ are the Lagrange multipliers, or the dual potentials.
The optimal $P^{*}$ is determined by the first-order optimality condition $\frac{\partial\mathcal{L}}{\partial P}|_{P^{*}}=0$, and can be expressed as a function of the dual potentials $f$ and $g$,
$P^{*}_{ij}:=a_{i}b_{j}\exp\left(\frac{f^{*}_{i}+g^{*}_{j}-c_{ij}}{\epsilon}\right).$ (18)
@@
$\sum_{j}P_{ij}=a_{i},\quad\sum_{i}P_{ij}=b_{j},$ (19)
and plugging the optimal $P^{*}_{ij}$ (18), we get
@@
If we solve the above marginal constraints for $\mathbf{f}^{*}$ and $\mathbf{g}^{*}$ by alternatively scaling $\mathbf{f}^{(k)}$ and $\mathbf{g}^{(k)}$ in the $\log$-domain, then we obtain the Sinkhorn algorithm with the LSE function.
@@ Appendix C Differentiation of EOT with respect to data
$\frac{d\mathrm{OT}_{\varepsilon}(\bm{\mu},\bm{\nu})}{d\mathbf{x}_{k}}=\frac{\partial\mathcal{L}(\mathbf{f},\mathbf{g},P)}{\partial\mathbf{x}_{k}}=\sum_{j=1}^{m}P_{kj}^{*}\frac{\partial c_{kj}}{\partial\mathbf{x}_{k}}=2\sum_{j=1}^{m}P_{kj}^{*}\big(\mathbf{x}_{k}-\mathbf{y}_{j}\big)=2a_{k}\Big(\mathbf{x}_{k}-T_{\varepsilon}(\mathbf{x}_{k})\Big),$ (21)
@@ E.1 Proof of Proposition 3
$P_{ij}:=P_{ij}(\hat{\mathbf{f}},\hat{\mathbf{g}})=a_{i}b_{j}\exp\!\left(\frac{\hat{f}_{i}+\hat{g}_{j}+(QK^{\top})_{ij}}{\varepsilon}\right).$
Recall $\bm{\delta}=\varepsilon\log\mathbf{b}$ and $\bm{\gamma}=\varepsilon\log\mathbf{a}$, so that $b_{j}=\exp(\delta_{j}/\varepsilon)$ and $a_{i}=\exp(\gamma_{i}/\varepsilon)$.
By definition,
$S_{X}(\hat{\mathbf{g}})_{ij}=\frac{(QK^{\top})_{ij}+\hat{g}_{j}+\delta_{j}}{\varepsilon},\qquad\exp\!\big(S_{X}(\hat{\mathbf{g}})_{ij}\big)=\exp\!\left(\frac{(QK^{\top})_{ij}+\hat{g}_{j}+\delta_{j}}{\varepsilon}\right).$
The row-update definition $)\hat{\mathbf{f}}^{+}=-\varepsilon\,\mathrm{LSE}_{\rm row}(S_{X}(\hat{\mathbf{g}}$)) means, entrywise,
$\hat{f}_{i}^{+}=-\varepsilon\log\sum_{j=1}^{m}\exp\!\big(S_{X}(\hat{\mathbf{g}})_{ij}\big)\quad\Longrightarrow\quad\sum_{j=1}^{m}\exp\!\big(S_{X}(\hat{\mathbf{g}})_{ij}\big)=\exp\!\left(-\frac{\hat{f}_{i}^{+}}{\varepsilon}\right).$
Now rewrite $P_{ij}$ as
$P_{ij}=a_{i}\exp\!\left(\frac{\hat{f}_{i}}{\varepsilon}\right)\;\exp\!\left(\frac{(QK^{\top})_{ij}+\hat{g}_{j}+\delta_{j}}{\varepsilon}\right)=a_{i}\exp\!\left(\frac{\hat{f}_{i}}{\varepsilon}\right)\exp\!\big(S_{X}(\hat{\mathbf{g}})_{ij}\big).$
Therefore the row-mass vector $r=P\mathbbm{1}_{m}$ satisfies
$r_{i}=\sum_{j=1}^{m}P_{ij}=a_{i}\exp\!\left(\frac{\hat{f}_{i}}{\varepsilon}\right)\sum_{j=1}^{m}\exp\!\big(S_{X}(\hat{\mathbf{g}})_{ij}\big)=a_{i}\exp\!\left(\frac{\hat{f}_{i}-\hat{f}_{i}^{+}}{\varepsilon}\right),$
which is the claimed expression for $r$. Moreover,
$\frac{P_{ij}}{r_{i}}=\frac{\exp\!\big(S_{X}(\hat{\mathbf{g}})_{ij}\big)}{\sum_{j^{\prime}=1}^{m}\exp\!\big(S_{X}(\hat{\mathbf{g}})_{ij^{\prime}}\big)}=\rm{Softmax}\!\big(S_{X}(\hat{\mathbf{g}})\big)_{ij},$
where $\rm{Softmax}(\cdot)$ is applied row-wise. Hence, in matrix form,
$),P=\mathrm{diag}(r)\,\rm{Softmax}\!\big(S_{X}(\hat{\mathbf{g}}$$\big)$,
and multiplying by any $V\in\mathbb{R}^{m\times p}$ gives
$PV=\mathrm{diag}(r)\,\rm{Softmax}\!\big(S_{X}(\hat{\mathbf{g}})\big)\,V.$
The column form $P^{\top}U=\mathrm{diag}(c)\,\rm{Softmax}\!\big(S_{Y}(\hat{\mathbf{f}})\big)\,U$ is shown similarly.
If $(\hat{\mathbf{f}}^{*},\hat{\mathbf{g}}^{*})$ satisfies the Sinkhorn fixed-point relations, then $\hat{\mathbf{f}}^{*}=\hat{\mathbf{f}}^{*+}$ and $\hat{\mathbf{g}}^{*}=\hat{\mathbf{g}}^{*+}$, so $r=\mathbf{a}\odot\exp((\hat{\mathbf{f}}^{*}-\hat{\mathbf{f}}^{*+})/\varepsilon)=\mathbf{a}$ and $c=\mathbf{b}\odot\exp((\hat{\mathbf{g}}^{*}-\hat{\mathbf{g}}^{*+})/\varepsilon)=\mathbf{b}$. Hence $P(\hat{\mathbf{f}}^{*},\hat{\mathbf{g}}^{*})$ has the prescribed marginals and coincides with the coupling induced by the optimal dual potentials, i.e., it recovers the optimal transport matrix. ∎
@@ G.1 Algorithms
A direct implementation of (10) would invoke separate kernels to (i) compute the score tiles $QK^{\top}$, (ii) add biases, and (iii) perform a row-wise $\mathrm{LSE}$ reduction, typically incurring extra HBM traffic and kernel-launch overhead due to intermediate score materialization. Instead, we fuse these stages into a single streaming kernel: each thread block loads a tile of $Q$ and iterates over tiles of $K$, updates the running row-wise max and rescaled sum-exp statistics $(\mathbf{m}_{I},\mathbf{s}_{I})$ (online $\mathrm{LSE}$), and emits $\hat{\mathbf{f}}_{I}$ (or $\hat{\mathbf{g}}_{I}$) once per row block. For transport matrix application, we fuse the corresponding softmax-weighted value accumulation into the same pass, analogous to fused attention kernels that compute $\rm{Softmax}(QK^{\top})V$ without storing $QK^{\top}$ or $\rm{Softmax}(QK^{\top})$ explicitly.

### v2

@@ Abstract
Entropic optimal transport (EOT) via Sinkhorn iterations is widely used in modern machine learning, yet GPU solvers remain inefficient at scale. Tensorized implementations suffer quadratic HBM traffic from dense $n\times m$ interactions, while existing online backends avoid storing dense matrices but still rely on generic tiled map-reduce reduction kernels with limited fusion. We present FlashSinkhorn, an IO-aware EOT solver for squared Euclidean cost that rewrites stabilized log-domain Sinkhorn updates as row-wise LogSumExp reductions of biased dot-product scores, the same normalization as transformer attention. This enables FlashAttention-style fusion and tiling: fused Triton kernels stream tiles through on-chip SRAM and update dual potentials in a single pass, substantially reducing HBM IO per iteration while retaining linear-memory operations. We further provide streaming kernels for transport application, enabling scalable first- and second-order optimization. On A100 GPUs, FlashSinkhorn achieves up to $32\times$ forward-pass and $161\times$ end-to-end speedups over state-of-the-art online baselines on point-cloud OT, improves scalability on OT-based downstream tasks. For reproducibility, we release an open-source implementation at https://github.com/ot-triton-lab/flash-sinkhorn.
@@ 2.1 EOT and Sinkhorn Algorithm
where $\Pi(\mathbf{a},\mathbf{b})=\{P\in\mathbb{R}_{\geq 0}^{n\times m}:P\mathbbm{1}_{m}=\mathbf{a},P^{\top}\mathbbm{1}_{n}=\mathbf{b}\}$ is the transport polytope, $C\in\mathbb{R}^{n\times m}$ is the cost matrix with entries $C_{ij}=\|\mathbf{x}_{i}-\mathbf{y}_{j}\|_{2}^{2}$, i.e., the squared Euclidean distance, and $\textrm{KL}\left(P\middle\|\mathbf{a}\otimes\mathbf{b}\right):=\sum_{i=1}^{n}\sum_{j=1}^{m}\left(P_{ij}\log\frac{P_{ij}}{a_{i}b_{j}}-P_{ij}+a_{i}b_{j}\right)$, denotes the relative entropy with $\varepsilon>0$ the regularization strength.
@@ 3.1 Streaming Sinkhorn Iterations
5:  Load $X_{I}$ to on-chip SRAM
@@
8:   Load $Y_{J}$, $\hat{\mathbf{g}}_{J}$ and $\bm{\delta}_{J}$ to on-chip SRAM
9:   $S\leftarrow(2X_{I}Y_{J}^{\top}+\hat{\mathbf{g}}_{J}+\bm{\delta}_{J})/\varepsilon$ // score tile + bias
@@ 3.2 Streaming Transport Matrix Application
5:  Load $X_{I}$, $\hat{\mathbf{f}}_{I}$, $\mathbf{a}_{I}$ to on-chip SRAM
6:  $\mathbf{m}_{I}\leftarrow$-$\infty$, $O_{I}\leftarrow 0$
@@
8:   Load $Y_{J}$, $\hat{\mathbf{g}}_{J}$, $\bm{\delta}_{J}$, $V_{J}$ to on-chip SRAM
9:   $S\leftarrow(2\,X_{I}Y_{J}^{\top}+\hat{\mathbf{g}}_{J}+\bm{\delta}_{J})/\varepsilon$ // score tile + target bias
@@
12:   $O_{I}\leftarrow e^{\mathbf{m}_{I}-\mathbf{m}_{\mathrm{new}}}\odot O_{I}+e^{S-\mathbf{m}_{\mathrm{new}}}\,V_{J}$ // online weighted sum
13:   $\mathbf{m}_{I}\leftarrow\mathbf{m}_{\mathrm{new}}$
14:  end for
15:  $\mathrm{out}_{I}\leftarrow\mathbf{a}_{I}\odot\exp\!\left(\hat{\mathbf{f}}_{I}/\varepsilon+\mathbf{m}_{I}\right)\odot O_{I}$ // source marginal correction
16:  Write $\mathrm{out}_{I}$ to HBM
17:end for
@@ Appendix B Derivation of Sinkhorn algorithm
where $\Pi(\mathbf{a},\mathbf{b})=\{P\in\mathbb{R}_{\geq 0}^{n\times m}:P\mathbbm{1}_{m}=\mathbf{a},P^{\top}\mathbbm{1}_{n}=\mathbf{b}\}$ , $C\in\mathbb{R}^{n\times m}$ is the cost matrix with entries $C_{ij}=\|\mathbf{x}_{i}-\mathbf{y}_{j}\|^{2}_{2}$, and $\mathrm{KL}(P\|\mathbf{a}\otimes\mathbf{b})=\sum_{ij}\left(P_{ij}\log\frac{P_{ij}}{a_{i}b_{j}}-P_{ij}+a_{i}b_{j}\right)$ is the Kullback–Leibler (KL) divergence (also called relative entropy) and $\varepsilon$ is the entropy regularization strength.
@@
$\mathcal{L}$ $(C,{\mathbf{f}},{\mathbf{g}},P)=\sum_{ij}\Big[C_{ij}P_{ij}+\varepsilon P_{ij}\log\frac{P_{ij}}{a_{i}b_{j}}-\varepsilon P_{ij}-f_{i}(P_{ij}-a_{i})-g_{j}(P_{ij}-b_{j})\Big]+\epsilon.$
where $\mathbf{f}$ and $\mathbf{g}$ are the Lagrange multipliers, or the dual potentials. The strong duality holds, then $\text{OT}_{\varepsilon}(\bm{\mu},\bm{\nu})=\max_{\mathbf{f},\mathbf{g}}\min_{P}\mathcal{L}(\mathbf{f},\mathbf{g},P)$.
The optimal $P^{*}$ is determined by the first-order optimality condition $\frac{\partial\mathcal{L}}{\partial P}=0$, and can be expressed as a function of the dual potentials $\mathbf{f}$ and $\mathbf{g}$,
$P_{ij}(\mathbf{f},\mathbf{g}):=a_{i}b_{j}\exp\left(\frac{f_{i}+g_{j}-C_{ij}}{\epsilon}\right).$ (18)
@@
$\sum_{j}P^{*}_{ij}=a_{i},\quad\sum_{i}P^{*}_{ij}=b_{j},$ (19)
The first-order optimality conditions of the outer problem yield,
@@
If we solve the above marginal constraints for $\mathbf{f}^{*}$ and $\mathbf{g}^{*}$ by alternatively scaling $\mathbf{f}^{(k)}$ and $\mathbf{g}^{(k)}$ in the $\log$-domain, then we obtain the Sinkhorn algorithm with the LSE function. Then the optimal transport matrix is
$P^{*}_{ij}=a_{i}b_{j}\exp\left(\frac{f^{*}_{i}+g^{*}_{j}-C_{ij}}{\epsilon}\right).$ (20)
@@ Appendix C Differentiation of EOT with respect to data
$\frac{d\mathrm{OT}_{\varepsilon}(\bm{\mu},\bm{\nu})}{d\mathbf{x}_{k}}=\frac{\partial\mathcal{L}(C,\mathbf{f},\mathbf{g},P^{*}(\mathbf{f},\mathbf{g}))}{\partial\mathbf{x}_{k}}=\sum_{j=1}^{m}P_{kj}^{*}\frac{\partial C_{kj}}{\partial\mathbf{x}_{k}}=2\sum_{j=1}^{m}P_{kj}^{*}\big(\mathbf{x}_{k}-\mathbf{y}_{j}\big)=2a_{k}\Big(\mathbf{x}_{k}-T_{\varepsilon}(\mathbf{x}_{k})\Big),$ (22)
@@ E.1 Proof of Proposition 3
$P_{ij}=a_{i}\exp\!\left(\frac{\hat{f}_{i}}{\varepsilon}\right)\exp\!\big(S_{X}(\hat{\mathbf{g}})_{ij}\big).$
For the matrix-vector product $(PV)_{i}=\sum_{j}P_{ij}V_{j}$, we have
$(PV)_{i}=a_{i}\exp\!\left(\frac{\hat{f}_{i}}{\varepsilon}\right)\sum_{j=1}^{m}\exp\!\big(S_{X}(\hat{\mathbf{g}})_{ij}\big)\,V_{j}.$
For numerical stability, let $m_{i}:=\max_{j}S_{X}(\hat{\mathbf{g}})_{ij}$ and factor out $\exp(m_{i})$:
$(PV)_{i}=a_{i}\exp\!\left(\frac{\hat{f}_{i}}{\varepsilon}+m_{i}\right)\sum_{j=1}^{m}\exp\!\big(S_{X}(\hat{\mathbf{g}})_{ij}-m_{i}\big)\,V_{j}.$ (26)
This formulation admits online (streaming) computation: as we iterate over column blocks $J$, we maintain a running maximum $m_{i}$ and accumulate $O_{i}:=\sum_{j}\exp(S_{ij}-m_{i})V_{j}$ using the recurrence
$O_{i}\leftarrow\exp(m_{i}^{\mathrm{old}}-m_{i}^{\mathrm{new}})\cdot O_{i}+\exp(S_{iJ}-m_{i}^{\mathrm{new}})\,V_{J}.$
After processing all blocks, the final output is $a_{i}\exp(\hat{f}_{i}/\varepsilon+m_{i})\cdot O_{i}$.
Define $s_{i}:=\sum_{j}\exp(S_{X}(\hat{\mathbf{g}})_{ij}-m_{i})$. Then
$\hat{f}_{i}^{+}=-\varepsilon\,\mathrm{LSE}_{j}(S_{X}(\hat{\mathbf{g}})_{ij})=-\varepsilon(m_{i}+\log s_{i}),$
and the row marginal satisfies
$r_{i}=a_{i}\exp\!\left(\frac{\hat{f}_{i}-\hat{f}_{i}^{+}}{\varepsilon}\right)=a_{i}\exp\!\left(\frac{\hat{f}_{i}}{\varepsilon}+m_{i}+\log s_{i}\right)=a_{i}\exp\!\left(\frac{\hat{f}_{i}}{\varepsilon}+m_{i}\right)\cdot s_{i}.$
Hence the formula eq. 15 equals
$(PV)_{i}=\frac{r_{i}}{s_{i}}\cdot O_{i}=r_{i}\cdot\frac{O_{i}}{s_{i}}=r_{i}\cdot\mathrm{Softmax}(S_{X}(\hat{\mathbf{g}}))_{i}\,V,$ (27)
confirming $PV=\mathrm{diag}(\mathbf{r})\,\mathrm{Softmax}(S_{X}(\hat{\mathbf{g}}))\,V$.
The streaming formulation eq. 26 is preferable for implementation because the normalizing sum $s_{i}$ cancels algebraically and need not be computed explicitly. This saves one accumulator per row and one reduction operation per tile.
The transpose case eq. 16 follows by symmetry, exchanging source and target roles.
At the Sinkhorn fixed point, $\hat{\mathbf{f}}^{*}=\hat{\mathbf{f}}^{*+}$ and $\hat{\mathbf{g}}^{*}=\hat{\mathbf{g}}^{*+}$, so $r_{i}=a_{i}$ and $c_{j}=b_{j}$. Thus $P(\hat{\mathbf{f}}^{*},\hat{\mathbf{g}}^{*})$ has the prescribed marginals $\mathbf{a}$ and $\mathbf{b}$, recovering the optimal transport matrix. ∎
@@ G.1 Algorithms
A direct implementation of (10) would invoke separate kernels to (i) compute the score tiles $QK^{\top}$, (ii) add biases, and (iii) perform a row-wise $\mathrm{LSE}$ reduction, typically incurring extra HBM traffic and kernel-launch overhead due to intermediate score materialization. Instead, we fuse these stages into a single streaming kernel: each thread block loads a tile of $Q$ and iterates over tiles of $K$, updates the running row-wise max and rescaled sum-exp statistics $(\mathbf{m}_{I},\mathbf{s}_{I})$ (online $\mathrm{LSE}$), and emits $\hat{\mathbf{f}}_{I}$ (or $\hat{\mathbf{g}}_{I}$) once per row block. For transport matrix application, we additionally maintain a weighted accumulator $O_{I}:=\sum_{j}\exp(S_{ij}-m_{i})\,V_{j}$ in the same streaming pass. Crucially, the normalizing sum $\mathbf{s}_{I}=\sum_{j}\exp(S_{ij}-m_{i})$ need not be tracked explicitly: it cancels algebraically in the final marginal correction $\mathrm{out}_{I}=\mathbf{a}_{I}\odot\exp(\hat{\mathbf{f}}_{I}/\varepsilon+\mathbf{m}_{I})\odot O_{I}$, saving one accumulator register per row and one reduction per tile compared to standard softmax-weighted attention.